# Bayesian Group Regularization in Generalized Linear Models with a Continuous Spike-and-Slab Prior thanks: Keywords and phrases: generalized linear model, group sparsity, spike-and-slab, spike-and-slab group lasso, posterior contraction, rate of convergence, variable selection

### v7

@@ Appendix A Proofs for Section 3
$\qquad\leq P\left(\exists~{}g\in S_{0}^{c},~{}~{}\lVert h_{g}(\bm{\beta}_{0})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
$\qquad\qquad+P\left(\exists~{}g\in S_{0}^{c},~{}~{}\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
$\qquad\leq\sum_{g=s_{0}+1}^{G}P\left(\lVert h_{g}(\bm{\beta}_{0})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
$\qquad\qquad+\sum_{g=s_{0}+1}^{G}P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
@@
$P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\lVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
$\qquad<P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\lVert_{2}\geq\frac{\lambda_{0}}{2}\right)$
$\qquad\leq P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})\rVert_{2,\infty}\lVert\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\geq\frac{\lambda_{0}}{2}\right)$
$\qquad\leq P\left(\lVert\mathbf{X}_{S_{0}}^{\top}\bm{\Sigma}(\widetilde{\bm{\beta}})\mathbf{X}_{g}\rVert_{2,\infty}\lVert\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\geq\frac{\lambda_{0}}{2}\right)$
$\qquad\lesssim\frac{2n}{\lambda_{0}},$ (A.24)
@@ B.1 Proof of Theorem 2
$\max\left\{K(f_{0i},f_{i}),V(f_{0i},f_{i})\right\}$
$\qquad\leq b^{\prime\prime}(\theta_{0i})(\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0}))^{2}(\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2}+o((\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2})$
$\qquad=(h^{-1})^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0})\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0})(\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2}+o((\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2}).$
Note that $h^{-1}$ and $\xi$ are strictly increasing. Thus, it must be the case that if $\max_{1\leq i\leq n}\{(h^{-1})^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0})\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0})(\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2}\}\leq\epsilon_{n}^{2}\rightarrow 0$, then $|\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0}|\rightarrow 0$ for all $1\leq i\leq n$. Recall $\bm{\Omega}(\bm{\beta}_{0})=\textrm{diag}\{(h^{-1})^{\prime}(\mathbf{X}\bm{\beta}_{0})\}\textrm{diag}\{\xi^{\prime}(\mathbf{X}\bm{\beta}_{0})\}$. Then both $n^{-1}\sum_{i=1}^{n}K(f_{0i},f_{i})$ and $n^{-1}\sum_{i=1}^{n}V(f_{0i},f_{i})$ can be bounded above by a constant multiple of $n^{-1}\lVert{\color[rgb]{0,0,0}\bm{\Omega}^{1/2}(\bm{\beta}_{0})}\mathbf{X}(\bm{\beta}-\bm{\beta}_{0})\rVert_{2}^{2}$ on $\mathcal{B}_{n}$.
@@
$\Pi(\mathcal{B}_{n})\geq$ $~{}\theta^{s_{0}}(1-\theta)^{G-s_{0}}e^{-\lambda_{1}\lVert\bm{\beta}_{S_{0}}\rVert_{2}}e^{-\lambda_{1}C_{2}/2\sqrt{n}}\left[1-\frac{4nGm_{\max}(m_{\max}+1)}{\lambda_{0}^{2}C_{2}^{2}s_{0}}\right]^{G-s_{0}}$
×$\qquad\times\prod_{g\in S_{0}}\frac{C_{g}}{m_{g}!}\left(\frac{\lambda_{1}C_{2}}{2s_{0}\sqrt{n}}\right)^{m_{g}}.$ (B.5)
@@
$-\log\Pi(\mathcal{B}_{n})$
$\qquad\lesssim 2cs_{0}\log G+\sqrt{s_{0}}\log G+\sum_{g\in S_{0}}\left[\log(m_{g}!)-\log(C_{g})+m_{g}\log\left(\frac{2s_{0}\sqrt{n}}{\lambda_{1}C_{2}}\right)\right].$ (B.6)
@@
$\sum_{S:|S|>K_{1}s_{0}}\theta^{|S|}\leq\sum_{k=\lfloor C_{1}s_{0}\rfloor+1}^{G}{G\choose k}\left(\frac{1}{G^{2}}\right)^{k}$ $\leq\sum_{k=\lfloor K_{1}s_{0}\rfloor+1}^{G}\left(\frac{e}{kG}\right)^{k}$
$<\sum_{k=\lfloor K_{1}s_{0}\rfloor+1}^{G}\left(\frac{e}{G\lfloor K_{1}s_{0}\rfloor+1}\right)^{k}$
$\leq G^{-(\lfloor K_{1}s_{0}\rfloor+1)}\lesssim e^{-K_{1}n\epsilon_{n}^{2}}.$ (B.9)
@@ B.2 Proof of Theorem 3
Note that $\pi(\bm{\beta}_{g})\geq\Psi_{1}(\bm{\beta}_{g}\mid\lambda_{0})$ for all $\bm{\beta}_{g}$. Further, for $\bm{\beta}_{g}\sim\Psi_{1}(\bm{\beta}_{g}\mid\lambda_{0})$, we have that $\lVert\bm{\beta}_{g}\rVert_{2}$ follows a gamma distribution with shape parameter $m_{g}$ and scale parameter $\lambda_{0}$. Thus,
@@
$H^{2}(f_{\bm{\beta}_{1},i},f_{\bm{\beta}_{2},i})$
$\qquad\leq\frac{K(f_{\bm{\beta}_{1},i},f_{\bm{\beta}_{2},i})}{2}$
$\qquad=\frac{(h^{-1})^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{1})\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{1})}{2}(\mathbf{x}_{i}^{\top}\bm{\beta}_{1}-\mathbf{x}_{i}^{\top}\bm{\beta}_{2})^{2}+o((\mathbf{x}_{i}^{\top}\bm{\beta}_{1}-\mathbf{x}_{i}^{\top}\bm{\beta}_{2})^{2}).$
@@
$\leq\frac{1}{2}\max_{1\leq g\leq G}\left(\lambda_{\max}\left(\frac{1}{n}\mathbf{X}_{g}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{1})}\mathbf{X}_{g}\right)\right)\left(\sum_{g=1}^{G}\lVert\bm{\beta}_{1g}-\bm{\beta}_{2g}\rVert_{2}\right)^{2}$
$\qquad+o\left(\lVert\mathbf{X}(\bm{\beta}_{1}-\bm{\beta}_{2})\rVert_{2}^{2}\right)$
@@
$N(\epsilon_{n}/36,\mathcal{A}_{n}^{\star},H_{n})$ $\leq N\left(\frac{C_{5}\epsilon_{n}}{36G},\mathcal{A}_{n}^{\star},\lVert\cdot\rVert_{2}\right)$
$=N\left(\frac{C_{5}\epsilon_{n}}{36G},\left\{\bm{\beta}\in\mathcal{A}_{n},\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\right\},\lVert\cdot\rVert_{2}\right).$ (B.16)
@@
$\{\bm{\beta}\in\mathcal{A}_{n},\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\}$
$\qquad\subset\left\{\bm{\beta}\in\mathbb{R}^{p}:\lVert\bm{\beta}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}},\textrm{ and }\lVert\bm{\beta}_{g}\rVert_{2}\leq\widetilde{\omega}\textrm{ for all }g\in S_{0}^{c}\right\}$
$\qquad\subset\left\{\lVert\bm{\beta}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\right\}\times[-\widetilde{\omega},\widetilde{\omega}]^{G-K_{1}s_{0}}.$
@@
(${G\choose K_{1}s_{0}}D\left(\frac{C_{5}\epsilon_{n}}{36G\sqrt{2}},\left\{\lVert\bm{\beta}_{S}-\bm{\beta}_{0S}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\right\},\lVert\cdot\rVert_{2}\right)$
($\qquad\leq{G\choose K_{1}s_{0}}\left(\frac{108\sqrt{2}G^{{\color[rgb]{0,0,0}M}+1}}{C_{5}\epsilon_{n}}\right)^{K_{1}s_{0}},$ (B.17)
@@
$\log N(\epsilon_{n}/36,\mathcal{A}_{n}^{\star},H_{n})$ $\leq K_{1}s_{0}\log G+K_{1}s_{0}\left[({\color[rgb]{0,0,0}M}+1)\log G+\log(108\sqrt{2n})\right]$
$\lesssim n\epsilon_{n}^{2},$ (B.18)
@@
$\qquad\leq\sum_{g\in\mathcal{S}_{0}}P(X_{g}>G^{{\color[rgb]{0,0,0}M}}-{\color[rgb]{0,0,0}C_{7}}\sqrt{G}\log G)$
$\qquad\leq\sum_{g\in S_{0}}\exp\left[-\lambda_{1}(G^{{\color[rgb]{0,0,0}M}}-{\color[rgb]{0,0,0}C_{7}}\sqrt{G}\log G)+m_{\max}\right]$
$\qquad=\exp\left[-\lambda_{1}G^{{\color[rgb]{0,0,0}M}}+\lambda_{1}{\color[rgb]{0,0,0}C_{7}}\sqrt{G}\log G+m_{\max}+\log s_{0}\right],$ (B.21)
@@
$\Pi(\mathcal{A}_{n}\setminus\mathcal{A}_{n}^{\star})$ $\leq 2\exp\left[-\lambda_{1}G^{{\color[rgb]{0,0,0}M}}+\lambda_{1}{\color[rgb]{0,0,0}C_{7}}\sqrt{G}\log G+\log G+m_{\max}\right]$
$\lesssim\exp(-C_{4}s_{0}\log G),$ (B.23)

### v8

@@ Appendix A Proofs for Section 3
$\qquad\leq P\left(\exists~{}g\in S_{0}^{c},~{}~{}\lVert h_{g}(\bm{\beta}_{0})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)+P\left(\exists~{}g\in S_{0}^{c},~{}~{}\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
$\qquad\leq\sum_{g=s_{0}+1}^{G}P\left(\lVert h_{g}(\bm{\beta}_{0})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)+\sum_{g=s_{0}+1}^{G}P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\rVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$
@@
$P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\lVert_{2}\geq\frac{\lambda_{g}^{\star}(\bm{0}_{m_{g}})}{2}\right)$ $<P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})^{\top}(\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}})\lVert_{2}\geq\frac{\lambda_{0}}{2}\right)$
$\leq P\left(\lVert\nabla_{S_{0}}h_{g}(\widetilde{\bm{\beta}})\rVert_{2,\infty}\lVert\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\geq\frac{\lambda_{0}}{2}\right)$
$\leq P\left(\lVert\mathbf{X}_{S_{0}}^{\top}\bm{\Sigma}(\widetilde{\bm{\beta}})\mathbf{X}_{g}\rVert_{2,\infty}\lVert\widehat{\bm{\beta}}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\geq\frac{\lambda_{0}}{2}\right)$
$\lesssim\frac{2n}{\lambda_{0}},$ (A.25)
@@ B.1 Proof of Theorem 2
$\max\left\{K(f_{0i},f_{i}),V(f_{0i},f_{i})\right\}$ $\leq b^{\prime\prime}(\theta_{0i})(\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0}))^{2}(\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2}+o((\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2})$
$=(h^{-1})^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0})\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0})(\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2}+o((\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2}).$
Note that $h^{-1}$ and $\xi$ are strictly increasing. Thus, if $\max_{1\leq i\leq n}\{(h^{-1})^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0})\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{0})(\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0})^{2}\}\leq\epsilon_{n}^{2}\rightarrow 0$, then $|\mathbf{x}_{i}^{\top}\bm{\beta}-\mathbf{x}_{i}^{\top}\bm{\beta}_{0}|\rightarrow 0$ for all $1\leq i\leq n$. Recall $\bm{\Omega}(\bm{\beta}_{0})=\textrm{diag}\{(h^{-1})^{\prime}(\mathbf{X}\bm{\beta}_{0})\}\textrm{diag}\{\xi^{\prime}(\mathbf{X}\bm{\beta}_{0})\}$. Then both $n^{-1}\sum_{i=1}^{n}K(f_{0i},f_{i})$ and $n^{-1}\sum_{i=1}^{n}V(f_{0i},f_{i})$ can be bounded above by a constant multiple of $n^{-1}\lVert{\color[rgb]{0,0,0}\bm{\Omega}^{1/2}(\bm{\beta}_{0})}\mathbf{X}(\bm{\beta}-\bm{\beta}_{0})\rVert_{2}^{2}$ on $\mathcal{B}_{n}$.
@@
$\Pi(\mathcal{B}_{n})\geq$ $~{}\theta^{s_{0}}(1-\theta)^{G-s_{0}}e^{-\lambda_{1}\lVert\bm{\beta}_{S_{0}}\rVert_{2}}e^{-\lambda_{1}C_{2}/2\sqrt{n}}\left[1-\frac{4nGm_{\max}(m_{\max}+1)}{\lambda_{0}^{2}C_{2}^{2}s_{0}}\right]^{G-s_{0}}\times\prod_{g\in S_{0}}\frac{C_{g}}{m_{g}!}\left(\frac{\lambda_{1}C_{2}}{2s_{0}\sqrt{n}}\right)^{m_{g}}.$ (B.5)
@@
$-\log\Pi(\mathcal{B}_{n})\lesssim 2cs_{0}\log G+\sqrt{s_{0}}\log G+\sum_{g\in S_{0}}\left[\log(m_{g}!)-\log(C_{g})+m_{g}\log\left(\frac{2s_{0}\sqrt{n}}{\lambda_{1}C_{2}}\right)\right].$ (B.6)
@@
$\sum_{S:|S|>K_{1}s_{0}}\theta^{|S|}\leq\sum_{k=\lfloor C_{1}s_{0}\rfloor+1}^{G}{G\choose k}\left(\frac{1}{G^{2}}\right)^{k}\leq\sum_{k=\lfloor K_{1}s_{0}\rfloor+1}^{G}\left(\frac{e}{kG}\right)^{k}$ $<\sum_{k=\lfloor K_{1}s_{0}\rfloor+1}^{G}\left(\frac{e}{G\lfloor K_{1}s_{0}\rfloor+1}\right)^{k}$
$\leq G^{-(\lfloor K_{1}s_{0}\rfloor+1)}$
$\lesssim e^{-K_{1}n\epsilon_{n}^{2}}.$ (B.9)
@@ B.2 Proof of Theorem 3
Note that $\pi(\bm{\beta}_{g})\geq\Psi(\bm{\beta}_{g}\mid\lambda_{0})$ for all $\bm{\beta}_{g}$. Further, for $\bm{\beta}_{g}\sim\Psi(\bm{\beta}_{g}\mid\lambda_{0})$, we have that $\lVert\bm{\beta}_{g}\rVert_{2}$ follows a gamma distribution with shape parameter $m_{g}$ and scale parameter $\lambda_{0}$. Thus,
@@
$H^{2}(f_{\bm{\beta}_{1},i},f_{\bm{\beta}_{2},i})$ $\leq\frac{K(f_{\bm{\beta}_{1},i},f_{\bm{\beta}_{2},i})}{2}=\frac{(h^{-1})^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{1})\xi^{\prime}(\mathbf{x}_{i}^{\top}\bm{\beta}_{1})}{2}(\mathbf{x}_{i}^{\top}\bm{\beta}_{1}-\mathbf{x}_{i}^{\top}\bm{\beta}_{2})^{2}+o((\mathbf{x}_{i}^{\top}\bm{\beta}_{1}-\mathbf{x}_{i}^{\top}\bm{\beta}_{2})^{2}).$
@@
$\leq\frac{1}{2}\max_{1\leq g\leq G}\left(\lambda_{\max}\left(\frac{1}{n}\mathbf{X}_{g}^{\top}{\color[rgb]{0,0,0}\bm{\Omega}(\bm{\beta}_{1})}\mathbf{X}_{g}\right)\right)\left(\sum_{g=1}^{G}\lVert\bm{\beta}_{1g}-\bm{\beta}_{2g}\rVert_{2}\right)^{2}+o\left(\lVert\mathbf{X}(\bm{\beta}_{1}-\bm{\beta}_{2})\rVert_{2}^{2}\right)$
@@
$N(\epsilon_{n}/36,\mathcal{A}_{n}^{\star},H_{n})$ $\leq N\left(\frac{C_{5}\epsilon_{n}}{36G},\mathcal{A}_{n}^{\star},\lVert\cdot\rVert_{2}\right)=N\left(\frac{C_{5}\epsilon_{n}}{36G},\left\{\bm{\beta}\in\mathcal{A}_{n},\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\right\},\lVert\cdot\rVert_{2}\right).$ (B.16)
@@
$\{\bm{\beta}\in\mathcal{A}_{n},\lVert\bm{\beta}-\bm{\beta}_{0}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\}$$\subset\left\{\bm{\beta}\in\mathbb{R}^{p}:\lVert\bm{\beta}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}},\textrm{ and }\lVert\bm{\beta}_{g}\rVert_{2}\leq\widetilde{\omega}\textrm{ for all }g\in S_{0}^{c}\right\}$
$\subset\left\{\lVert\bm{\beta}_{S_{0}}-\bm{\beta}_{0S_{0}}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\right\}\times[-\widetilde{\omega},\widetilde{\omega}]^{G-K_{1}s_{0}}.$
@@
(${G\choose K_{1}s_{0}}D\left(\frac{C_{5}\epsilon_{n}}{36G\sqrt{2}},\left\{\lVert\bm{\beta}_{S}-\bm{\beta}_{0S}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\right\},\lVert\cdot\rVert_{2}\right)\leq{G\choose K_{1}s_{0}}\left(\frac{108\sqrt{2}G^{{\color[rgb]{0,0,0}M}+1}}{C_{5}\epsilon_{n}}\right)^{K_{1}s_{0}},$ (B.17)
@@
$\log N(\epsilon_{n}/36,\mathcal{A}_{n}^{\star},H_{n})$ $\leq K_{1}s_{0}\log G+K_{1}s_{0}\left[({\color[rgb]{0,0,0}M}+1)\log G+\log(108\sqrt{2n})\right]\lesssim n\epsilon_{n}^{2},$ (B.18)
@@
$\leq\sum_{g\in\mathcal{S}_{0}}P(X_{g}>G^{{\color[rgb]{0,0,0}M}}-{\color[rgb]{0,0,0}C_{7}}\sqrt{G}\log G)$
$\leq\sum_{g\in S_{0}}\exp\left[-\lambda_{1}(G^{{\color[rgb]{0,0,0}M}}-{\color[rgb]{0,0,0}C_{7}}\sqrt{G}\log G)+m_{\max}\right]$
$=\exp\left[-\lambda_{1}G^{{\color[rgb]{0,0,0}M}}+\lambda_{1}{\color[rgb]{0,0,0}C_{7}}\sqrt{G}\log G+m_{\max}+\log s_{0}\right],$ (B.21)
@@
$\Pi(\mathcal{A}_{n}\setminus\mathcal{A}_{n}^{\star})$ $\leq 2\exp\left[-\lambda_{1}G^{{\color[rgb]{0,0,0}M}}+\lambda_{1}{\color[rgb]{0,0,0}C_{7}}\sqrt{G}\log G+\log G+m_{\max}\right]\lesssim\exp(-C_{4}s_{0}\log G),$ (B.23)